# Hyperbolic Random Graphs: Clique Number and Degeneracy with Implications for Colouring

### v1

@@ 1 Introduction
Many real-world networks have a heterogeneous degree distribution, close to a power-law, as well as a constant clustering coefficient. The hyperbolic random graph model (HRG) introduced by Krioukov et. al. [27] combines both properties [18], which has led to considerable interest in recent year. Various aspects of HRGs have been studied, including clique size [7, 8, 15], treewidth [9], minimum vertex cover size [3, 4], and diameter [16, 23, 31]. A hyperbolic random graph is a graph embedded in the hyperbolic plane where pairs of vertices have an edge if they are close according to the hyperbolic distance. It is generated by randomly throwing $n$ vertices on a disk of radius $R$ (dependent on $n$). Since hyperbolic space grows exponentially, most vertices are of small degree and located close to the boundary of the disk, while few vertices have large degree and lie close to the centre. This distribution of vertices leads to the power-law degree distribution of HRGs.
@@
We aim to study these structural parameters that are not only fundamental to the model but also in general for algorithm design in various models of computation [1, 11, 17]. The most prominent large clique in an HRG is formed by the vertices in the graph’s core [7]. Simply put, the core emerges among the polynomially many vertices of distance at most $R/2$ from the centre of the disk, which due to the triangle inequality form a clique. At this point one may wonder whether the core forms the largest clique of the graph, a statement that we disprove in Proposition 4.7. Nevertheless we show that the largest clique can at most be a small constant factor larger than the core.
Theorem (Informal version of Theorem 4.10).
There exists a constant $\delta>0$ such that for any threshold HRG $G$, $\sigma(G)+1\leq\omega(G)\leq\sqrt{2-\delta}\cdot\sigma(G)$ holds w.e.h.p.  [^1] An event holds with extremely high probability (w.e.h.p.), if for if for every $c>1$, there exists an $n_{0}$ such that for every $n\geq n_{0}$ the event holds with probability at least $1-n^{-c}$.
@@
We can now see that the core and clique are not the same. Nonetheless, (i) this theorem shows that the largest clique size and the core size are closely related, and (ii) the core of HRGs is a very well understood object, whereas the largest clique is not. Thus the natural next step is to bound the degeneracy using the core. We show the following theorem.
Theorem (Informal version of Theorems 3.5 and 3.5).
There exists constants $\delta_{1},\delta_{2}>0$ such that for any threshold HRG $G$, $(1+\delta_{1})\sigma(G)\leq\kappa(G)\leq(2-\delta_{2})\cdot\sigma(G)$ holds w.e.h.p.
The main surprise of this theorem is that the degeneracy is bounded away from the coresize by a constant factor. As the chromatic number is lower bounded by the core size, the upper bound on the degeneracy in this theorem implies a simple algorithm colouring with at most $(2-\delta_{2})\chi(G)$ colours. The approximation guarantee of this algorithm ranges from $2/\sqrt{3}$ to $4/3$ depending on further model parameters, see Section 2 and Theorem 3.8 for details. In any case, this improves on the previously best approximation ratio of $2$ [6, Lemma 7]. As discussed earlier this algorithm iteratively removes vertices of degree at most $(2-\delta_{2})\sigma(G)-1$ and then colours them in the reverse order. The next thing one would hope is to be able colour the graph with $\omega(G)$ colours using the same process. In [2], the authors conducted experiments where they were iteratively removing the vertex with the smallest degree of the graph, up to vertices with residual degree equal to the $\omega(G)$. In their findings, this process did not remove every vertex, implying $\omega(G)$ < $\kappa(G)$ for their generated graphs. We substantiate their findings by providing a rigorous proof demonstrating that the clique number is, in fact, a constant factor smaller than the degeneracy.
Theorem (Informal version of Theorem 4.3).
@@
Our final contribution is to study the degeneracy of geometric inhomogeneous random graphs (GIRGs) [22], a sibling to HRGs. The GIRGs also combine heterogeneity and high clustering. For most properties GIRGs and HRGs exhibit the same behavior. Perhaps the first paper to find a difference between them is [8], where the authors show that the minimum number of maximal cliques in the two models is not the same. We show that the degeneracy between HRGs and GIRGs is signifcantly different, see Figure 1 and Corollary 5.5.
Outline. See Figure 1 for a table with our results, as well as a plot comparing the bounds of our theorems for various model parameters. In Section 1.1 we provide a detailed discussion of our results and techniques. Section 3 contains bounds on the degeneracy of HRGs (Theorem 3.5). In Section 4, we show the gap between clique number and degeneracy (Theorem 4.3), as well as bounds on the clique number (Theorem 4.10). Finally, Section 5 contains results about the degeneracy of GIRGs. Statements where proofs or details are omitted due to space constraints can be found in the provided appendix.
Figure 1: Results on the degeneracy $\kappa(G)$ and the size of the largest clique $\omega(G)$ in hyperbolic random graphs (HRG) and geometric inhomogeneous random graphs (GIRG). The bounds hold w.e.h.p. and are stated in comparison to the core size $\sigma(G)$. The dependency on the model parameter $\alpha\in(1/2,1)$ is plotted on the x-axis with respect to the multiplicative factor in front of the core size $\sigma(G)$ on the y-axis. Prior work is listed with white background, whereas our results are listed with blue background.
@@ 1.1 Discussion of our Results and Techniques
As discussed HRGs have a power-law degree distribution [18, 32], that is, the probability that a vertex has degree $k$ is given by $\sim k^{-(2\alpha+1)}$ The model parameter $\alpha\in(1/2,1)$ controls the power-law exponent and all our results, particularly the size of the aforementioned constant factor gap depends on the choice of $\alpha$. For the easy of presentation this overview largely omits this dependence, but the summary of our results in Figure 1 plots it in detail.
@@
Figure 2: Illustration of the inner-neighbourhood. (a) The grey area is the ball $\mathcal{B}_{0}(r)$. The hatched area $\mathcal{I}(r)=\mathcal{B}_{u}(R)\cap\mathcal{B}_{0}(r)$ is the inner-ball of $u$. The vertices in $\mathcal{I}(r)$ form the inner-neighbourhood of $u$. (b) Sketch of proof of Lemma 3.2. Vertex $u^{*}$ is the vertex with the largest inner-degree. Each vertex $v\in U=V\cap\mathcal{B}_{0}(r^{*})$ has at least nearly as many neighbours within $\mathcal{B}_{0}(r^{*})$ as $|\operatorname{\Gamma}(u^{*})|$ (w.e.h.p.)
@@
Lower bound on degeneracy (Theorem 3.5). Surprisingly, we show in Lemma 3.2 that the maximum inner-degree also produces a lower bound on the degeneracy $\kappa(G)$, in the sense that $\kappa(G)\geq(1-o(1))|\operatorname{\Gamma}(u^{*})|$ w.e.h.p. This yields asymptotically tight bounds on $\kappa(G)$. We prove the lower bound of $\kappa(G)$ by considering the subgraph $G^{\prime}$ induced by the vertices that have smaller radius than $u^{*}$ (see Figure 2 b). Note that $G^{\prime}$ contains vertices that are not neighbours of $u^{*}$. We show that every vertex $u$ in $G^{\prime}$ has, w.e.h.p., at least $(1-o(1))|\operatorname{\Gamma}(u^{*})|$ neighbours in $G^{\prime}$; this statement uses the choice of $u^{*}$ and is not true if $u$ was an arbitrary vertex of the graph. Now, in any ordering of the vertices of $G^{\prime}$, the first vertex has at least $(1-o(1))|\operatorname{\Gamma}(u^{*})|$ neighbours of greater index, implying $(1-o(1))|\operatorname{\Gamma}(u^{*})|\leq\kappa(G^{\prime})\leq\kappa(G)$. While this provides a lower bound that asymptotically matches our upper bound, a lower order gap remains.
@@
Gap between clique number and degeneracy (Theorem 4.3). The most immediate lower bound for degeneracy is the clique number [1], because in every ordering of the vertices of $G$, the vertex of the clique with the lowest index has at least $\omega(G)-1$ neighbours of higher index. We prove that the lower bounds on the degeneracy that are derived from the clique number are strictly worse than the bounds discussed above obtained via analysing the inner-degree. Our approach to show this is the following: we take an arbitrary clique $K$ and the vertex $u$ with the largest radius in $K$. Let $G^{\prime\prime}$ be the subgraph induced by $\operatorname{\Gamma}(u)$. Next, we partition $G^{\prime\prime}$ intro three sets of vertices, each containing at least a constant fraction of the vertices of $G^{\prime\prime}$ w.e.h.p. See Figure 3 b for an illustration of this partition. Lastly, we use purely geometric arguments to show that $K$ cannot contain vertices from all three sets. As the left out set contains a constant fraction of $u$’s inner neighbourhood, implying the gap. This gap between $\omega(G)$ and $\kappa(G)$ and makes approaches for computing the clique number via degeneracy, like that of Walteros and Buchanan [33], unsuitable for HRGs.
@@
Clique number and the core (Theorem 4.10). The upper bound on the clique number is proven via a geometric approach. Let $u,v,w$ be any three vertices. If the vertices are far apart they cannot be contained in a clique. Otherwise their pairwise distance is at most $R$. Using the hyperbolic version Jung’s theorem [19, 13, 14] implies that they are contained in a ball $\mathcal{B}$ of small radius and we show that $\mathcal{B}$ also has a small area. Hence the expected number of vertices in $\mathcal{B}$ is small as well. As this expectation is well concentrated whenever the area is significantly large, this bound holds with extremely high probability when introducing lower order deviations from the expectation. Via a union bound over all possible $\binom{n}{3}$ triples of vertices we rule out that any clique contained in such a covering ball is large. The main claim now follows because any clique has to be contained in one of these coverings balls.
The question of whether $\omega(G)/\sigma(G)\to 1$ as $n\to\infty$ remains open.
@@ 2 Preliminaries
We follow the formalisation of hyperbolic random graphs introduced in [32], which is known as the native representation. We denote by $\mathbb{H}^{2}=[0,\infty)\times[0,2\pi)$ the hyperbolic plane in the polar coordinate system, where points $x\in\mathbb{H}^{2}$ is parameterised by a radius $r(x)$ and an angle $\varphi(x)$. We equip $\mathbb{H}^{2}$ with a metric $\operatorname{d_{h}}(x,y)$ characterised by
@@
The geometric space of most importance in this work is the bounded disk in $\mathbb{H}^{2}$ defined by $\mathcal{D}_{R}=[0,R]\times[0,2\pi)$, where $R=2\log(n)+C$ with $C\in\Theta(1)$. We refer to point $(0,0)$ as the centre of this disk. The space $\mathcal{D}_{R}$ inherits the topology of $\mathbb{H}^{2}$, and from now on we shall only consider subsets of this space – thus, for example, a ball around a point $y\in\mathcal{D}_{R}$ is defined by the set $\mathcal{B}_{\varepsilon}(y)=\{x\in\mathcal{D}_{R}:\operatorname{d_{h}}(x,y)\leq\varepsilon\}\subseteq\mathcal{D}_{R}$.
@@
where $\mathrm{d}x$ denotes the Lebesgue measure on $\mathcal{D}_{R}$. This measure differs from the uniform probability measure on $\mathcal{D}_{R}$ in that it puts more mass at the centre of the disk; both measures coincide at $\alpha=1$. The benefit of $\mu$ lies in the properties it induces in our central object of study, the hyperbolic random graph.
@@
A (threshold) hyperbolic random graph or HRG is a pair $G=(V,E)$ defined by the following procedure. First, $n$ vertices are sampled independently at random in $\mathcal{D}_{R}$ according to $\mu$. Then any two vertices $u,v\in V$ are connected by an edge if and only if their distance $\operatorname{d_{h}}(u,v)$ is at most $R$. We write $G\sim\mathcal{G}(n,\alpha,C)$ to denote a graph generated in this way.
@@
We briefly note that a variant of HRGs exists in which vertices are not connected purely according to whether their distance is below a threshold, but rather with probability $p(u,v)=(1+\exp\left(\frac{1}{2T}\left(\operatorname{d_{h}}(u,v)-R\right)\right))^{-1}$ determined by both distance and a “temperature” parameter $T$ (see e.g. [28, §3.1]).
Degeneracy, clique number, chromatic number and core. For a graph $G=(V,E)$, the degeneracy $\kappa(G)$ is the minimum integer $k$ for which there exists an ordering of the vertex set of $G$, $V=(v_{1},v_{2},\cdots,v_{n})$, such that for every index $i\in[n-1]$, $v_{i}$ has at most $k$ neighbours with greater index. The clique number $\omega(G)$ is the size of the largest clique of $G$. The chromatic number $\chi(G)$ is the smallest number of colours required so that a conflict-free vertex colouring is possible for $G$. The core of a hyperbolic random graph is the set of vertices with radius at most $R/2$. Since for any points $u,v\in\mathcal{B}_{0}(R/2)$ the distance is at most $\operatorname{d_{h}}(u,v)\leq R$, the core forms a clique. We denote the size of the core by $\sigma(G)$. Finally, we note that the following chain of inequalities holds: $\sigma(G)\leq\omega(G)\leq\chi(G)\leq\kappa(G)+1.$

### v2

@@ 1 Introduction
Many real-world networks have a heterogeneous degree distribution, close to a power-law, as well as a constant clustering coefficient. The hyperbolic random graph model (HRG) introduced by Krioukov et. al. [28] combines both properties [18], which has led to considerable interest in recent years. Various aspects of HRGs have been studied, including clique size [7, 8, 15], treewidth [9], minimum vertex cover size [3, 4], and diameter [16, 23, 32]. A hyperbolic random graph is a graph embedded in the hyperbolic plane where pairs of vertices have an edge if they are close according to the hyperbolic distance. It is generated by randomly throwing $n$ vertices on a disk of radius $R$ (dependent on $n$). Since hyperbolic space grows exponentially, most vertices are of small degree and located close to the boundary of the disk, while few vertices have large degree and lie close to the centre. This distribution of vertices leads to the power-law degree distribution of HRGs.
@@
We aim to study these structural parameters that are not only fundamental to the model but also in general for algorithm design in various models of computation [1, 11, 17]. The most prominent large clique in an HRG is formed by the vertices in the graph’s core [7]. Simply put, the core emerges among polynomially many vertices of distance at most $R/2$ from the centre of the disk, which due to the triangle inequality form a clique. We denote the size of this clique by $\sigma(G)$. At this point one may wonder whether the core forms the largest clique of the graph, a statement that we disprove in Proposition 4.8. Nevertheless we show that the largest clique can at most be small constant factor larger than the core.
Theorem (Simplified version of Theorem 4.13).
There exists a constant $\delta>0$ such that for any threshold HRG $G$, $\sigma(G)+1\leq\omega(G)\leq\sqrt{4/3-\delta}\cdot\sigma(G)$ holds w.e.h.p.  [^1] An event holds with extremely high probability (w.e.h.p.), if for every $c>1$, there exists an $n_{0}$ such that for every $n\geq n_{0}$ the event holds with probability at least $1-n^{-c}$.
@@
We can now see that the core and clique are not the same. Nonetheless, (i) this theorem shows that the largest clique size and the core size are closely related, and (ii) the core of HRGs is a very well understood object, whereas the largest clique is not. Thus the natural approach to bound the degeneracy is to use the core. We show the following theorem.
Theorem (Simplified version of Theorems 3.7 and 3.7).
There exist constants $\delta_{1},\delta_{2}>0$ such that for any threshold HRG $G$, $(1+\delta_{1})\cdot\sigma(G)\leq\kappa(G)\leq(4/3-\delta_{2})\cdot\sigma(G)$ holds w.e.h.p.
The main surprise of this theorem is that the degeneracy is bounded away from the coresize by a constant factor. As the chromatic number is lower bounded by the core size, the upper bound on the degeneracy in this theorem implies a simple algorithm colouring with at most $(4/3-\delta_{2})\chi(G)$ colours. The approximation guarantee of this algorithm ranges from $2/\sqrt{3}$ to $4/3$ depending on further model parameters, see Section 2 and Theorem 3.10 for details. In any case, this improves on the previously best approximation ratio of $2$ [6, Lemma 7]. The algorithm iteratively removes vertices of degree at most $(4/3-\delta_{2})\sigma(G)-1$ and then colours them in the reverse order. The next thing one would hope is to be able colour the graph with $\omega(G)$ colours using the same process. In [2], the authors conducted experiments where they were iteratively removing the vertex with the smallest degree of the graph, up to vertices with residual degree equal to $\omega(G)$. In their findings, this process did not remove every vertex, implying $\omega(G)$ < $\kappa(G)$ for their generated graphs. We substantiate their findings by providing a rigorous proof demonstrating that the clique number is, in fact, a constant factor smaller than the degeneracy.
Theorem (Simplified version of Theorem 4.3).
@@
Our final contribution is to study the degeneracy of geometric inhomogeneous random graphs (GIRGs) [22], a sibling to HRGs. The GIRGs also combine heterogeneity and high clustering. For most properties GIRGs and HRGs exhibit the same behaviour. Perhaps the first paper to find a difference between them is [8], where the authors show that the minimum number of maximal cliques in the two models differ. We show a significant discrepancy for the degeneracy of GIRGs compared to that of HRGs, see Figure 1 and Corollary 5.5.
Outline. See Figure 1 for a table with our results, as well as a plot comparing the bounds of our theorems for various model parameters. In Section 1.1 we provide a detailed discussion of our results and techniques. Section 3 contains bounds on the degeneracy of HRGs (Theorem 3.7). In Section 4, we show the gap between clique number and degeneracy (Theorem 4.3), as well as bounds on the clique number (Theorem 4.13). Finally, Section 5 contains results about the degeneracy of GIRGs. Statements where proofs or technical details are omitted can be found in the Appendix A.
Figure 1: Results on the degeneracy $\kappa(G)$ and the size of the largest clique $\omega(G)$ in hyperbolic random graphs (HRG) and geometric inhomogeneous random graphs (GIRG). The bounds hold w.e.h.p. and are stated in comparison to the core size $\sigma(G)$. Each curve represents the multiplicative factor in front of $\sigma(G)$ for $\kappa(G)$ and $\omega(G)$ (y-axis) depending on the parameter $\alpha\in(1/2,1)$ (x-axis). Prior work is listed with white background, whereas our results are listed with blue background.
@@ 1.1 Discussion of our Results and Techniques
HRGs have a power-law degree distribution [18, 33], that is, the probability that a vertex has degree $k$ is given by $\sim k^{-(2\alpha+1)}$ The model parameter $\alpha\in(1/2,1)$ controls the power-law exponent and all our results, particularly the size of the aforementioned constant factor gap depends on the choice of $\alpha$. For the ease of presentation this overview largely omits this dependence, but the summary of our results in Figure 1 plots it in detail.
@@
Figure 2: Illustration of the inner-neighbourhood. (a) The pink area is the ball $\mathcal{B}_{0}(r)$. The hatched area $\mathcal{I}_{u}=\mathcal{B}_{u}(R)\cap\mathcal{B}_{0}(r)$ is the inner-ball of $u$. The vertices $V\cap\mathcal{I}_{u}$ form the inner-neighbourhood $\operatorname{\Gamma}(u)$. (b) Sketch of proof of Lemma 3.2. Vertex $u^{*}$ is the vertex with the largest inner-degree. Each vertex $v\in U=V\cap\mathcal{B}_{0}(r^{*})$ has at least nearly as many neighbours within $\mathcal{B}_{0}(r^{*})$ as $|\operatorname{\Gamma}(u^{*})|$ (w.e.h.p.)
@@
Lower bound on degeneracy (Theorem 3.7). In Lemma 3.2 we show that the maximum inner-degree also produces a lower bound on the degeneracy $\kappa(G)$, in the sense that $\kappa(G)\geq(1-o(1))|\operatorname{\Gamma}(u^{*})|$ w.e.h.p. This yields asymptotically tight bounds on $\kappa(G)$. We prove the lower bound of $\kappa(G)$ by considering the subgraph $G^{\prime}$ induced by the vertices that have smaller radius than $u^{*}$ (see Figure 2 b). Note that $G^{\prime}$ contains vertices that are not neighbours of $u^{*}$. We show that every vertex $u$ in $G^{\prime}$ has, w.e.h.p., at least $(1-o(1))|\operatorname{\Gamma}(u^{*})|$ neighbours in $G^{\prime}$; this statement uses the choice of $u^{*}$ and is not true if $u$ was an arbitrary vertex. Now, in any ordering of the vertices of $G^{\prime}$, the first vertex has at least $(1-o(1))|\operatorname{\Gamma}(u^{*})|$ neighbours of greater index, implying $(1-o(1))|\operatorname{\Gamma}(u^{*})|\leq\kappa(G^{\prime})\leq\kappa(G)$. While this provides a lower bound that asymptotically matches our upper bound, a lower order gap remains.
@@
Gap between clique number and degeneracy (Theorem 4.3). The most immediate lower bound for degeneracy is the clique number [1], because in every ordering of the vertices of $G$, the vertex of the clique with the lowest index has at least $\omega(G)-1$ neighbours of higher index. We prove that the lower bounds on the degeneracy that are derived from the clique number are strictly worse than the bounds discussed above obtained via analysing the inner-degree. Our approach to show this is the following: we take an arbitrary clique $K$ and the vertex $u$ with the largest radius in $K$. Let $G^{\prime\prime}$ be the subgraph induced by $\operatorname{\Gamma}(u)$. Next, we partition $G^{\prime\prime}$ intro three sets of vertices, each containing at least a constant fraction of the vertices of $G^{\prime\prime}$ w.e.h.p. See Figure 3 b for an illustration of this partition. Lastly, we use purely geometric arguments to show that $K$ cannot contain vertices from all three sets. The gap then follows as the left out set contains a constant fraction of $u$’s inner neighbourhood. This gap between $\omega(G)$ and $\kappa(G)$ makes approaches for computing the clique number via degeneracy, like that of Walteros and Buchanan [34], unsuitable for HRGs.
@@
Clique number and the core (Theorem 4.13). The upper bound on the clique number is proven via a geometric approach. Let $u,v,w$ be any three vertices. If the vertices are far apart they cannot be contained in a clique. Otherwise their pairwise distance is at most $R$. Using the hyperbolic of version Jung’s theorem [19, 13, 14] implies that they are contained in a ball $\mathcal{B}$ of small radius and we show that $\mathcal{B}$ also has a small area. Hence the expected number of vertices in $\mathcal{B}$ is small as well. As this expectation is well concentrated whenever the area is significantly large, this bound holds with extremely high probability when introducing lower order deviations from the expectation. Via a union bound over all possible $\binom{n}{3}$ triples of vertices we rule out that any clique contained in such a covering ball is large. The main claim now follows because any clique has to be contained in one of these coverings balls. The question of whether $\omega(G)/\sigma(G)\to 1$ as $n\to\infty$ remains open.
@@ 2 Preliminaries
We follow the formalisation of hyperbolic random graphs introduced in [33], which is known as the native representation. We denote by $\mathbb{H}^{2}=[0,\infty)\times[0,2\pi)$ the hyperbolic plane in the polar coordinate system, where a point $x\in\mathbb{H}^{2}$ is parametrised by a radius $r(x)$ and an angle $\varphi(x)$. We equip $\mathbb{H}^{2}$ with a metric $\operatorname{d_{h}}(x,y)$ characterised by
@@
The geometric space of most importance in this work is the bounded disk in $\mathbb{H}^{2}$ defined by $\mathcal{D}_{R}=[0,R]\times[0,2\pi)$, where $R=2\log(n)+C$ with $C\in\Theta(1)$. We refer to point $(0,0)$ as the centre of this disk. The space $\mathcal{D}_{R}$ inherits the topology of $\mathbb{H}^{2}$, and from now on we shall only consider subsets of this space – thus, for example, a ball around a point $x\in\mathcal{D}_{R}$ is defined by the set $\mathcal{B}_{x}(\varepsilon)=\{y\in\mathcal{D}_{R}:\operatorname{d_{h}}(x,y)\leq\varepsilon\}\subseteq\mathcal{D}_{R}$.
@@
where $\rho$ is the density of $\mu$ with respect to the Lebesgue measure on $\mathcal{D}_{R}$. This measure differs from the uniform probability measure on $\mathcal{D}_{R}$ in that it puts more mass at the centre of the disk; both measures coincide at $\alpha=1$. The benefit of $\mu$ lies in the properties it induces in our central object of study, the hyperbolic random graph.
@@
A (threshold) hyperbolic random graph or HRG is a pair $G=(V,E)$ defined by the following procedure. First, $n$ vertices are sampled independently at random in $\mathcal{D}_{R}$ according to $\mu$. Then any two vertices $u,v\in V$ are connected by an edge if and only if their distance $\operatorname{d_{h}}(u,v)$ is at most $R$. We write $G\sim\mathcal{G}(n,\alpha,C)$ to denote a graph generated in this way. A vertex $u\in V$ is identified by its point coordinates in $\mathbb{H}^{2}$ and we write $V\cap\mathcal{A}$ to denote the set of vertices that are located in an area $\mathcal{A}\subseteq\mathcal{D}_{R}$.
@@
We briefly note that a variant of HRGs exists in which vertices are not connected purely according to whether their distance is below a threshold, but rather with probability $p(u,v)=(1+\exp\left(\frac{1}{2T}\left(\operatorname{d_{h}}(u,v)-R\right)\right))^{-1}$ determined by both distance and a “temperature” parameter $T$ (see e.g. [29, §3.1]).
Degeneracy, clique number, chromatic number and core. For a graph $G=(V,E)$, the degeneracy $\kappa(G)$ is the minimum integer $k$ for which there exists an ordering of the vertex set of $G$, $V=(v_{1},v_{2},\cdots,v_{n})$, such that for every index $i\in[n-1]$, $v_{i}$ has at most $k$ neighbours with greater index. The clique number $\omega(G)$ is the size of the largest clique of $G$. The chromatic number $\chi(G)$ is the smallest number of colours required so that a conflict-free vertex colouring is possible for $G$. The core of a hyperbolic random graph is the set of vertices with radius at most $R/2$ and we denote its size by $\sigma(G)$. Since for any points $u,v\in\mathcal{B}_{0}(R/2)$ the distance is at most $\operatorname{d_{h}}(u,v)\leq R$, the core forms a clique. Finally, since the core is a clique, any vertex of a clique needs a different colour in a conflict-free colouring, and $\chi(G)\leq\kappa(G)+1$ (see e.g. [30, Lemma 4]), we have the following chain of inequalities. {observation} Let $G\sim\mathcal{G}(n,\alpha,C)$ be a threshold HRG. Then,
$\sigma(G)\leq\omega(G)\leq\chi(G)\leq\kappa(G)+1.$
Concentration bounds. We use the following Chernoff bounds [31, Theorem 4.4].
Theorem 2.1 (Chernoff bound).
$\mathbb{P}\left(X\geq(1+\varepsilon)\mathbb{E}[X]\right)\leq e^{-\varepsilon^{2}\cdot\mathbb{E}[X]/3}\text{ and }\mathbb{P}\left(X\leq(1-\varepsilon)\mathbb{E}[X]\right)\leq e^{-\varepsilon^{2}\cdot\mathbb{E}[X]/2}.$